OpenStudy (anonymous):

For which system of equations is (–2, 3) the solution? A. 3x + 2y = 0 4x + 3y = –1 B. –3x + 4y = –6 2x + 5y = 11 C. –x + 7y = 16 5x – 3y = 16 D. –x + 7y = 23 5x + 3y = –1

OpenStudy (shamil98):

Plug them, into the equations. 3(-2) + 2(3) = 0 that's correct 4(-2) + 3(3) = -1 that's wrong so option A is wrong

OpenStudy (shamil98):

Let's try option b. -3(-2) + 4(3) = -6 6 + 12 = -6 Nope, not option b.

OpenStudy (anonymous):

uh huh

OpenStudy (shamil98):

Let's try option C. -(-2) + 7(3) = 16 4 + 21 = 16 nope..

OpenStudy (shamil98):

Through process of elimination, you know that option D is your correct answer.

OpenStudy (anonymous):

Lol thank you!